OpenStudy (anonymous):

How do I take the derivative of (integral(e^-s^2 ds) from 0 to t?

OpenStudy (jamesj):

Do you know the fundamental theorem of calculus?

OpenStudy (anonymous):

would it simply be e^(-s^2)?

OpenStudy (jamesj):

Yes indeed.

OpenStudy (anonymous):

Okay good thanks just reviewing

OpenStudy (anonymous):

but wait you don't take 0 or t into account?

OpenStudy (anonymous):

the integration from 0 to t I meant

OpenStudy (jamesj):

No. The Theorem says (minus some technicalities) the following. Let F(x) be the function \[F(x) = \int\limits_{a}^{x} f(t) \ dt\] for some function f. Then the derivative of F exists and \[\frac{dF}{dx}(x) = f(x)\] Notice that this result is completely independent of a.

OpenStudy (mrsmarsh):

You need to use the second fundamental theorem of calculus